ID |
Date |
Author |
Type |
Category |
Subject |
Status |
Text |
 |
279
|
Mon Feb 10 14:57:56 2020 |
Ino + Gael + Thierry | Other | Hardware | Azimuth Locking System | Fixed | Last week we put in sevice the new ALS
completly modified:
_ Main System with new reducer. |
14x |
278
|
Thu Jan 16 22:06:39 2020 |
tions applied to solve this issue :Martin Will, Thierry Le Flour | Telescope Operations | Software | Semi-Expert Drive GUI not connecting / MOS server not working | Fixed | In the late afternoon shifters noticed
the semi-expert drive GUI was not updating,
several restarts did not help. Checking into |
 |
277
|
Sat Dec 7 19:16:01 2019 |
Daniel Mazin, Koji Noda | Hardware | Hardware | ALS East: problem and preliminary solution | | On Dec 6, Dario, Frederic and Yuji showed
to Koji Noda and Daniel Mazin how to operate
the new ALS. The idea was to do an unlock |
6x |
276
|
Wed Nov 27 21:27:29 2019 |
Mathieu David Armand | Telescope Operations | | CDM & SG images with alternative target | | |
|
275
|
Tue Nov 26 21:37:03 2019 |
David Sanchez, Armand Fiasson, Mathieu De Bony | Telescope Operations | | CDM bending model image 26-11-2019 | | The aim was to observe several stars with
different magnitude and check the sensitivity
of the CDM. We wanted to use a semi-automatic |
|
274
|
Mon Nov 25 22:53:22 2019 |
Mathieu David Armand | Telescope Operations | | CDM images 19/11/25 | | Mg=1.25
DenebDark 60x1s
RA-2
TestTracking 60x1s Displacement
from Deneb RA-2
Deneb 60x1s
DenebLed |
|
273
|
Sun Nov 24 19:33:56 2019 |
Mathieu David Armand | Telescope Operations | General | CDM images for bending model | | - CDM images at park out position.
Exposure 100ms.
OARL on. |
|
272
|
Thu Sep 19 19:10:39 2019 |
Martin, Ino, Thierry, Quentin, Thomas | Tracking Debugging | | Star images with various Az and Zd offsets | | We tracked stars at various zenith angles
and changed the Az and Zd offsets, taking
images with the PSF camera for each position. |
|
271
|
Mon Sep 16 08:59:41 2019 |
Drive Team- Ino&Thierry + Ievgen + Vitalii + Jean-Luc | PLC + Control Software debugging | | Test TCU<->Drive | | Thierry&Ino on site
Ievgen, Vitalli and Jean-Luc remotely.
TCU tests to perform the parkin-in |
|
270
|
Thu Sep 12 16:41:53 2019 |
Drive Team- Ino&Thierry | PLC + Control Software debugging | | | | Tuesday September 110th :
Tests park-in and park-out
to manage the different type of failures
Tests |
|
269
|
Sat Jul 27 11:11:08 2019 |
Yusuke, Satoshi, Tomo, Yoshiki | Other | | GUI stacked due to a wrong shutdown procedure | | Drive GUI stacked due to a wrong shutdown
procedure (Jul. 27th 01-02UTC). This was
simply a human error. |
|
268
|
Tue Jul 23 13:50:43 2019 |
Yusuke, Patricia | Problem Fixed | | Reboot Flywheel PLC | | Regarding the problem on the Flywheel PLC
(see
here), we entered UC1 under the supervision of Daniel |
|
267
|
Mon Jul 22 23:07:20 2019 |
Yusuke, Daniel K., Mitsunari, Satoshi, Tomo, Juan Abel, Yoshiki | Telescope Operations | | Flywheel PLC down | | After finishing the normal start-up procedure,
we tried to park out by GUI, but got an error
(22:30 local time): "MSG_WARNINGPark_Out_TimeOut" |
 |
266
|
Fri Jul 19 23:06:13 2019 |
Mitsunari, Tomohiro, Satoshi, Moritz, Yusuke | Telescope Operations | Issues | Telescope stuck in strange state/lost control by Semi-Expert GUI | | Tonight, we tracked Vega to test (a correction
of) the pointing offset of the Drive.
Around 22:30UTC, we tried different |
|
265
|
Tue Jul 9 00:46:52 2019 |
Mitsunari, Yukiho, Noah, Moritz | Telescope Operations | Issues | Park-in problems | | At parking in, the telescope got stuck
at 94.254°. We checked in the container
that fuse Q21 went off. We did not see any |
|
264
|
Sat Jul 6 23:17:54 2019 |
Moritz, Dirk | Telescope Operations | | Park-in problems | | When parking in the telescope, accidently
hit "Shutdown" button instead of
"Parkin" in Drive Semi-epert GUI. |
|
263
|
Sat Jul 6 22:47:38 2019 |
Dirk Hoffmann | Telescope Operations | General | Car flash from road above+ | | When checking the telescope visually outside,
while it was in the "park-out"
position, a car happened to pass by from |
|
262
|
Fri Jul 5 23:23:57 2019 |
Dirk, Moritz | Telescope Operations | General | Park-in report | | On park-in command, telescope went to the
right position 95/-1, but did not recognise
that it was there. |
|
261
|
Fri Jul 5 08:58:42 2019 |
Dirk Hoffmann | Problem Fixed | Software | Small problem with park-in | | When parking in yesterday, the telescope
drove to the nominal position (approx. 95°
zenith angle and -1° azimuth), but then |
|
260
|
Wed Jul 3 21:02:09 2019 |
Dirk, Taka, Mitsunari, Yukiho, Noah | Telescope Operations | Issues | Feedback from shiftcrew | | (For Thierry with love <3 :-)
The semi-expert drive control panel
uses various terms for the zenith angle: |
|